find the mean, median, and mode of the set: 1,2,2,3,3,3,4,4,4,4
Answer by venugopalramana(3286)

find the mean, median, and mode of the set: 1,2,2,3,3,3,4,4,4,4
mean =(1+2*2+3*3+4*4)/10=30/10=3
mode is 4 as it has maximum frequency
median is the average of 5 th. and 6 th. numbers in the orderly sequence of numbers (which is same as that given above)that is 3